# WoVoGen: World Volume-aware Diffusion for Controllable Multi-camera Driving Scene Generation

### v1

@@ Abstract
Figure 1: Our WoVoGen is crafted to predict the ambient environment and generate reasonable visual feedback in response to the driving maneuvers of the ego vehicle. To unleash the power of generative models that come in leaps and bounds, we encode structured traffic information into a regular grid framework, i.e. world volume and devise a novel latent diffusion-based world model to autoregressively perform world volume prediction.
@@ 1 Introduction
Recent driving scene content synthesis techniques can be categorized into two main groups: rendering-based [41, 44, 9, 42] and diffusion-based [33, 43, 8, 39, 18] methods. Rendering-based methods benefit from an explicit 3D or 4D world structure, ensuring stringent 3D consistency. Yet, this approach tends to offer limited diversity and requires considerable effort to produce multi-camera videos that comply with real-world lighting, weather conditions, etc. On the other hand, methods based on finetuned stable diffusion models boast high diversity and can easily generate images that follow real-world distributions. Despite the noted advantages, diffusion-based approaches [33, 43, 8, 18, 39] that employ bounding boxes or HD maps as control signals offer a sparse representation of the driving scene. Consequently, the generated videos often lack critical information about the scene, leading to an incomplete understanding. This limitation underscores the necessity for a diffusion-based model that incorporates a dense representation of the world as a condition, ensuring a more comprehensive and accurate depiction of the driving environment.
@@ 2 Related work
Diffusion model for visual content generation
Diffusion models, as demonstrated in [11, 32, 4, 27], exhibit the capability to generate a diverse array of images through a learned denoising process. Building upon the foundation of a well-trained image generation model, several works have emerged to augment its controllable generation capabilities, such as text-guided approaches [5, 26, 29, 28], as well as layout-guided approaches [45, 14, 46, 19], etc. Further advancements in image generation extend to video generation, emphasizing temporal consistency in recent works [31, 47, 40, 1]. Additionally, there has been a surge in interest in novel view generation techniques [22, 30, 21, 24], focused on generating diverse perspectives from existing images. These advancements necessitate models with enhanced capabilities to ensure spatial consistency in generated outputs.
@@
Rendering-based driving scene synthesis
The rendering-based approaches [42, 41, 44, 9] initially reconstruct comprehensive world structures, followed by image or video rendering based on this reconstructed data. While these methods yield high-quality synthetic results, they often lack robust control capabilities. Efforts like NeuralField-LDM [16] have attempted to introduce diffusion models into volume rendering to address diversity issues and improve control in rendering-based synthesis. However, they still rely heavily on numerous real-world samples from various perspectives or additional supervision during model training, raising concerns about their efficacy.
Diffusion-based driving scene synthesis
The diffusion-based approaches [33, 43, 8, 39, 18] leverage pre-trained diffusion models and incorporate sparse conditions, i.e., HD maps and bounding boxes, enabling precise control over driving scene content generation. Part of the works [33, 43, 8] primarily focus on single-frame image generation. As a unified pipeline, they encode ground truth condition information and inject conditional features into the diffusion model to generate images corresponding to the specified condition. Other works [39, 18] have ventured into driving scene video generation. Building upon trained single-frame generation models, they finetune the temporal module to ensure the temporal coherence of generated elements across consecutive frames in the video.
@@ 3.2 Overall architecture
The proposed architecture, referred to as WoVoGen, comprises two distinct operational branches: the world model branch and the world volume-aware generation branch. The world model branch is responsible for generating future world volumes, incorporating action inputs and several initial frames of the world volume to inform its predictions. Meanwhile, the world volume-aware generation branch focuses on the generation of multi-camera video outputs based on temporal world volumes. An overview of our comprehensive pipeline is visually represented in Figure 2.
@@ 3.4 World model
Given a clip of multi-camera videos captured during driving, denoted as $\{\mathcal{I}_{0},\cdots,\mathcal{I}_{N_{\text{past}}}\}$, we initially employ off-the-shelf scene understanding models to infer the environment. This allows us to obtain the 3D world volume at these moments, represented as $\left\{\mathcal{W}_{t}\right\}_{t=1,\cdots,N_{\text{past}}}$. Subsequently, the world volume for the time instance $N_{\text{past}}+1$ is inferred, utilizing the past world volumes and the corresponding driving actions.
Figure 3: (a): an action attention block enhances the model by incorporating essential action information. (b): a guidance attention block integrates the CLIP feature of a specific object into the latent representation, further refining the model’s capabilities.
Latent world volume autoencoder
Given the world volume $\mathcal{W}\in\mathbb{R}^{Z\times H\times W\times C}$, the encoder is designed to compress $\mathcal{W}$ into its latent representation, $z_{w}=E_{\mathcal{W}}(\mathcal{W})\in\mathbb{R}^{\frac{Z}{s}\times\frac{H}{s}\times\frac{W}{s}\times C_{z}}$. Here, $s$ represents the downsampling factor, and is set equal to $Z$ to achieve a 2D latent representation. The decoder, $D_{\mathcal{W}}$, is then employed to reconstruct the original world volume from the latent representation $z_{w}$. The autoencoder is trained using a reconstruction loss and vq-regularization [6], aimed at minimizing the variance within the latent space.
Latent world diffusion
After the training of the latent autoencoder, we commence the training of the world model, with both the encoder and the decoder remaining frozen. As illustrated in Figure 2, our world model is a diffusion model that captures the conditional distribution of the latent world volume in the future, conditioned on the past world volumes and the driving actions of the ego vehicle. The noise predictor $\epsilon_{\phi}$ in this stage is implemented as a time-conditioned UNet [11]. The world volumes of preceding frames are first encoded by the same latent encoder to dimensions commensurate with the noised latent $z_{i}$, and then channel-wise concatenate with $z_{i}$ to serve as the input for the noise predictor.
@@
Jointly modeling of the consecutive frames
To enhance the temporal consistency of predicted future world volumes and achieve more realistic scene dynamics, we transition to generating consecutive world volumes by considering a joint distribution of a world volume sequence, $p(\mathcal{W}_{N_{\text{past}}+1},\cdots,\mathcal{W}_{N_{\text{past}}+N_{\text{future}}}|\mathcal{W}_{1},\cdots,\mathcal{W}_{N_{\text{past}}};\mathcal{A})$.
@@ 3.5 World volume-aware 2D feature
World volume encoding
The world volumes inherently possess semantic information, initially represented in the form of simplistic labels. To enhance their informativeness, we employ a featurization process utilizing CLIP [25]:
@@
Camera volume sampling
To incorporate the world volume into image generation, we sample from it using dense rays emitted from the camera. We construct a 3D grid for each camera’s frustum, denoted as $p_{c}$, with dimensions $D_{c}\times H_{c}\times W_{c}$. This grid is based on the camera’s intrinsic and extrinsic properties. The process is formalized as follows:
@@
where $\mathcal{F}_{c}\in\mathbb{R}^{B\times N_{\text{future}}\times C\times D_{c}\times H_{c}\times W_{c}}$ represents the camera frustum. We then apply a squeeze-and-excitation operation [12] on the depth channel and sum along the depth dimension to obtain the world volume-aware 2D image feature:
@@ 3.6 World volume-aware diffusion generation
Panoptic diffusion
We aggregate the world volume-aware 2D image feature from different view into a single panoptic feature $\mathcal{F}_{pano}$ input to the diffusion model:
@@
Scene guidance
Except for the conditional feature, we also introduce text prompt-based scene guidance into the latent representations. This involves extracting the CLIP feature from a text-based description of an image and mapping this text feature to the intermediate layers of the latent diffusion model, following a similar approach as LDM [27].
Object guidance
When it comes to objects that require specific placement within the generated image, we emphasize their pixel-level locations in the latent space using a cross-attention calculation.
@@ 3.7 Video generation
Temporal consistency
The temporal attention block is added to ensure the generation of multi-frame images consistent, which calculates the attention by:
@@
3.8 Training target
With all the conditions serving as inputs, the training objective can be formulated as:
$\min_{\theta}\mathcal{L}=\mathbb{E}_{\mathbf{z},\epsilon\sim\mathcal{N}(0,1),\tau}\left[\|\epsilon-\epsilon_{\theta}(\mathbb{\mathbf{z}}_{\tau},\tau,c)\|^{2}_{2}\right],$ (8)
where $\theta$ is the learnable parameters in our network, $c$ denotes the conditions.
@@ 4.1 Experimental setup
Dataset
We conducted experiments using the NuScenes [2] dataset which comprises 700 training videos and 150 validation videos. Our study utilized images generated by the six camera views trained on the NuScenes dataset.
For ground-truth occupancy data, we utilized CVPR 2023 3D Occupancy Prediction Challenge [35, 34], aligned with semantic labels consistent with the nuScenes-lidarseg dataset [7]. The ego car-centric local map is derived from the global HD map using the vehicle’s ego pose and precisely aligned with the occupancy data.
Data preprocessing
@@
Figure 6: Examples of controlled editing on nuScenes [2] validation dataset. Top: the ability to selectively add or remove specific objects (such as trees, buildings, cars, etc., highlighted by red circles in the figure) within the world volume empowers the precise and coherent generation of diverse driving scenarios across multiple cameras. Bottom: due to WoVoGen’s advanced 3D understanding capabilities, the rotation of perspectives across multiple cameras can be achieved by modifying the camera’s extrinsic parameters. This functionality enables the generation of driving scenes from any desired angle.
Training
The training process utilized eight NVIDIA A6000 GPUs, encompassing 30,000 iterations for the world volume auto-encoder, 50,000 iterations for volume diffusion, 50,000 iterations for single-frame generation training, and 3,500 iterations for refining video generation.
Inferring
We begin with three initial world volumes organized from ground-truth data and proceed to generate the subsequent three world volumes by incorporating actions as input. This iterative process continues, producing a quasi-long world volume video sequence. Ultimately, the world volume video sequence is decoded to form a multi-view camera video.
Evaluation
We conducted a comparative analysis of our method’s image quality against other approximate works in the following text. To quantitatively assess our work, we employed established metrics such as Fréchet Inception Distance [10] (FID) and Fréchet Video Distance [36] (FVD) for evaluating the quality of the generated video. Additionally, we provided qualitative demonstrations that underscore the advantages of our proposed method.
@@ 4.2 Results
4.2.1 World volume generation
@@ 4.2.2 Multi-camera single-frame image generation
Figure 5 shows the qualitative results of the multi-camera single-frame image generation under diverse conditions. We can observe that the generated images have highly consistency, which clearly shows that our model indeed learns to understand the geometric and photometric relationships between different views of the same object or scene. Furthermore, from the first row of Figure 5, we can see that the generated images accurately correspond to the occupancy guide. Beyond that, we can also control the style of the generated images through natural language. In the last two rows of the Figure 5, our WoVoGen demonstrates a profound comprehension of the specified textual conditions, skillfully creating highly realistic simulations of rainy day scenarios. Remarkably, this model also exhibits the capability to interpret and respond to location-specific instructions through natural language processing. When prompted with “Drive in Singapore Onenorth,” the model skillfully generates a scene embodying the garden city essence typical of that region. This notable feature not only illustrates the model’s advanced understanding of geographical nuances but also highlights its potential to generate a wide array of diverse and contextually accurate data.
@@ 4.2.3 Multi-camera single-frame image editing
Rearranging objects:
We can add, delete, or transform the objects within the occupancy. As shown in Figure 6, we can add trees, vehicles, or rearrange the location of objects.
Camera extrinsic editing:
More interestingly, we can rotate the camera stereo when projecting the world volume to the camera volume and generate the corresponding scene under a new camera setting.
Method Multi-view Multi-frame FIDnormal-$\downarrow$ FVDnormal-$\downarrow$
@@ 5 Conclusion
Appendix A Masks for object guidance
@@
Appendix B Additional results
B.1 Downstream tasks

### v4

# https://github.com/fudan-zvg/WoVoGen

@@ Abstract
Keywords:
Diffusion Autonomous driving World volume
Figure 1: Our WoVoGen is crafted to generate future world volumes (i.e., HD maps and occupancy) and high-quality multi-camera street-view images, with the input of past world-volumes. The bottom row shows the weather-based control of WoVoGen. Specifically, with the predicted world volume at time $t_{4}$ and a weather description, the multi-camera images transit from rainy to sunny conditions, while maintaining the street layout.
@@ 1 Introduction
Recent driving scene content synthesis techniques can be categorized into two main groups: rendering-based [43, 46, 10, 44, 3] and diffusion-based [34, 45, 9, 40, 18] methods. Rendering-based methods benefit from an explicit 3D or 4D world structure, ensuring stringent 3D consistency. Yet, this approach tends to offer limited diversity and requires considerable effort to produce multi-camera videos that comply with real-world lighting, weather conditions, etc. On the other hand, methods based on finetuned stable diffusion models boast high diversity and can easily generate images that follow real-world distributions. Despite the noted advantages, diffusion-based approaches [34, 45, 9, 18, 40] that employ bounding boxes or HD maps as control signals offer a sparse representation of the driving scene. Consequently, the generated videos often lack critical information about the scene, leading to an incomplete understanding. This limitation underscores the necessity for a diffusion-based model that incorporates a dense representation of the world as a condition, ensuring a more comprehensive and accurate depiction of the driving environment.
@@ 2 Related work
Diffusion model for visual content generation Diffusion models, as demonstrated in [12, 32, 5, 27], exhibit the capability to generate a diverse array of images through a learned denoising process. Building upon the foundation of a well-trained image generation model, several works have emerged to augment its controllable generation capabilities, such as text-guided approaches [6, 26, 29, 28], as well as layout-guided approaches [47, 15, 48, 19], etc. Further advancements in image generation extend to video generation, emphasizing temporal consistency in recent works [31, 49, 42, 1]. Additionally, there has been a surge in interest in novel view generation techniques [22, 30, 21, 24], focused on generating diverse perspectives from existing images. These advancements necessitate models with enhanced capabilities to ensure spatial consistency in generated outputs.
@@
Rendering-based driving scene synthesis The rendering-based approaches [44, 43, 46, 10] initially reconstruct comprehensive world structures, followed by image or video rendering based on this reconstructed data. While these methods yield high-quality synthetic results, they often lack robust control capabilities. Efforts like NeuralField-LDM [16] have attempted to introduce diffusion models into volume rendering to address diversity issues and improve control in rendering-based synthesis. However, they still rely heavily on numerous real-world samples from various perspectives or additional supervision during model training, raising concerns about their efficacy.
Diffusion-based driving scene synthesis The diffusion-based approaches [34, 45, 9, 40, 18] leverage pre-trained diffusion models and incorporate sparse conditions, i.e., HD maps and bounding boxes, enabling precise control over driving scene content generation. Part of the works [34, 45, 9] primarily focus on single-frame image generation. As a unified pipeline, they encode ground truth condition information and inject conditional features into the diffusion model to generate images corresponding to the specified condition. Other works [40, 18] have ventured into driving scene video generation. Building upon trained single-frame generation models, they finetune the temporal module to ensure the temporal coherence of generated elements across consecutive frames in the video.
Figure 3: (a): an action attention block enhances the model by incorporating action information. (b): a guidance attention block integrates the CLIP feature of a specific object into the latent representation.
@@ 3.2 Overall architecture
The proposed architecture, referred to as WoVoGen, comprises two distinct operational branches: the world model branch and the world volume-aware generation branch. The world model branch is responsible for generating future world volumes, incorporating action inputs and several initial frames of the world volume to inform its predictions. Meanwhile, the world volume-aware generation branch focuses on the generation of multi-camera video outputs based on temporal world volumes. An overview of our comprehensive pipeline is visually represented in Figure 2.xi
@@ 3.4 World model
Given a clip of multi-camera videos captured during driving, denoted as
$\{\mathcal{I}_{0},\cdots,\mathcal{I}_{N_{\text{past}}}\}$, we initially employ off-the-shelf scene understanding models BEVDet-occ [14] to infer the environment. This allows us to obtain the 3D world volume at these moments, represented as $\left\{\mathcal{W}_{t}\right\}_{t=1,\cdots,N_{\text{past}}}$. Subsequently, the world volume for the time instance $N_{\text{past}}+1$ is inferred, utilizing the past world volumes and the corresponding driving actions.
Latent world volume autoencoder Given the world volume $\mathcal{W}\in\mathbb{R}^{Z\times H\times W\times C}$, the encoder is designed to compress $\mathcal{W}$ into its latent representation, $z_{w}=E_{\mathcal{W}}(\mathcal{W})\in\mathbb{R}^{\frac{Z}{s}\times\frac{H}{s}\times\frac{W}{s}\times C_{z}}$. Here, $s$ represents the downsampling factor, and is set equal to $Z$ to achieve a 2D latent representation. The decoder, $D_{\mathcal{W}}$, is then employed to reconstruct the original world volume from the latent representation $z_{w}$. The autoencoder is trained using a reconstruction loss and vq-regularization [7], aimed at minimizing the variance within the latent space.
4D Latent world diffusion After the training of the latent autoencoder, we commence the training of the world model, with both the encoder and the decoder remaining frozen. As illustrated in Figure 2, our world model is a diffusion model that captures the conditional distribution of the latent world volume in the future, conditioned on the past world volumes and the driving actions of the ego vehicle. The noise predictor $\epsilon_{\phi}$ in this stage is implemented as a time-conditioned UNet [12]. The world volumes of preceding frames are first encoded by the same latent encoder to dimensions commensurate with the noised latent $z_{i}$, and then channel-wise concatenate with $z_{i}$ to serve as the input for the noise predictor.
@@
Jointly modeling of the consecutive frames To enhance the temporal consistency of predicted future world volumes and achieve more realistic scene dynamics, we transition to generating consecutive world volumes by considering a joint distribution of a world volume sequence.
@@ 3.5 World volume-aware 2D feature
World volume encoding The world volumes inherently possess semantic information, initially represented in the form of simplistic labels. To enhance their informativeness, we employ a featurization process utilizing CLIP [25]:
@@
Camera volume sampling To incorporate the world volume into image generation, we sample from it using dense rays emitted from the camera. We construct a 3D grid for each camera’s frustum, denoted as $p_{c}$, with dimensions $D_{c}\times H_{c}\times W_{c}$. This grid is based on the camera’s intrinsic and extrinsic properties. The process is formalized as follows:
@@
where $\mathcal{F}_{cam}\in\mathbb{R}^{B\times N_{\text{future}}\times C\times D_{c}\times H_{c}\times W_{c}}$ represents the camera frustum. We then apply a squeeze-and-excitation operation [13] on the depth channel and sum along the depth dimension to obtain the world volume-aware 2D image feature:
@@ 3.6 World volume-aware diffusion generation
Panoptic diffusion We aggregate the world volume-aware 2D image feature from different view into a single panoptic feature $\mathcal{F}_{pano}$ input to the diffusion model:
@@
Scene guidance Except for the conditional feature, we also introduce text prompt-based scene guidance into the latent representations. This involves extracting the CLIP feature from a text-based description of an image and mapping this text feature to the intermediate layers of the latent diffusion model, following a similar approach as LDM [27].
Object guidance When it comes to objects that require specific placement within the generated image, we emphasize their pixel-level locations in the latent space using a cross-attention calculation.
@@ 3.7 Video generation
Temporal consistency The temporal attention block is added to ensure the generation of multi-frame images consistent, which calculates the attention by:
@@
Figure 6: Examples of controlled editing on nuScenes [2] validation dataset. Top: the ability to selectively add or remove specific objects (such as trees, buildings, cars, etc., highlighted by red circles in the figure) within the world volume empowers the precise and coherent generation of diverse driving scenarios across multiple cameras. Bottom: due to WoVoGen’s advanced 3D understanding capabilities, the rotation of perspectives across multiple cameras can be achieved by modifying the camera’s extrinsic parameters.
@@ 4.1 Experimental setup
Dataset We conducted experiments using the nuScenes [2] dataset which comprises 700 training videos and 150 validation videos. Our study utilized images from the six camera views.
For ground truth occupancy, we use the data from CVPR 2023 3D Occupancy Prediction Challenge [36, 35], aligned with semantic labels consistent with the nuScenes-lidarseg dataset [8]. The ego car-centric local map is derived from the global HD map using the vehicle’s ego pose and precisely aligned with the occupancy data.
@@
Training The training process utilized eight NVIDIA A6000 GPUs, encompassing 30,000 iterations for the world volume auto-encoder, 50,000 iterations for volume diffusion, 50,000 iterations for single-frame generation training, and 3,500 iterations for refining video generation.
Inferring We begin with three initial world volumes organized from ground-truth data and proceed to generate the subsequent three world volumes by incorporating actions as input. This iterative process continues, producing a quasi-long world volume video sequence. Ultimately, the world volume video sequence is decoded to form a multi-view camera video.
Evaluation We conducted a comparative analysis of our method’s image quality against other approximate works in the following text. To quantitatively assess our work, we employed established metrics such as Fréchet Inception Distance [11] (FID) and Fréchet Video Distance [37] (FVD) for evaluating the quality of the generated video. Additionally, we provided qualitative demonstrations that underscore the advantages of our proposed method.
@@ 4.2 Results
4.2.1 4D World volume generation
@@ 4.2.2 Multi-camera single-frame image generation
Figure 5 shows the qualitative results of the multi-camera single-frame image generation under diverse conditions. We can observe that the generated images have highly consistency. Furthermore, from the first row of Figure 5, we can see that the generated images accurately correspond to the occupancy guide. Beyond that, we can also control the style of the generated images through natural language. In the last two rows of the Figure 5, our WoVoGen demonstrates a profound comprehension of the specified textual conditions, skillfully creating highly realistic simulations of rainy day scenarios. When prompted with “Drive in Singapore Onenorth,” the model skillfully generates a scene embodying the garden city essence typical of that region.
@@ 4.2.3 Multi-camera single-frame image editing
Rearranging objects: We can add, delete, or transform the objects within the occupancy. As shown in Figure 6, we can add trees, vehicles, or rearrange the location of objects.
Camera extrinsic editing: More interestingly, we can rotate the camera stereo when projecting the world volume to the camera volume and generate the corresponding scene under a new camera setting.
[…]
@@ 4.2.4 Multi-camera video generation
4.3 Ablation studies
4.3.1 Effectiveness of world volume and object guidance
In this study, we examine the significance of world volume and object guidance, as detailed in Table 2. We analyze three scenarios: object guidance alone, world volume alone, and the combination of world volume with object guidance. When solely relying on object guidance, interpreted as object mask information, the generated results are notably inferior. In contrast, incorporating world volume, regardless of the presence of object guidance, significantly enhances the generation outcomes. This underscores the critical importance of world volume in the generative process. Additionally, the incorporation of object guidance leads to more precise object localization, especially when evaluating the fidelity of the generated images, as indicated by the highest FID scores.
4.3.2 Consistency
To assess the quantitative outcomes of multi-camera and temporal consistency, we adopt a quantitative methodology utilizing Key Points Matching (KPM) [41]. This approach involves a pre-trained matching model [33] to calculate the average number of matching keypoints. Table 2 reveals that there is an impressive match rate of up to 80.2% for keypoints across successive frames, underscoring a significant accomplishment in maintaining temporal consistency. Regarding multi-view consistency, we observe that 90.0% of keypoints match between adjacent cameras, further highlighting our methodology’s efficacy in ensuring robust multi-camera consistency.
Table 2: Left: Quantitative comparison of image generation quality on the nuScenes validation set with and without object guidance. Right: Quantitative results of multi-view consistency and tmerpoal consistency. KPM is used for evaluation
[…]
$\times$ $\checkmark$ 28.6 Multi-view 90.0
$\checkmark$ $\times$ 27.8 Temporal 80.2
$\checkmark$ $\checkmark$ 27.6
@@ 5 Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62106050 and 62376060), Natural Science Foundation of Shanghai (Grant No. 22ZR1407500), USyd-Fudan BISA Flagship Research Program and Lingang Laboratory (Grant No. LG-QS-202202-07).
Appendix 0.A Appendix
0.A.1 Masks for object guidance
@@
0.A.2 Zero-shot generation
In Figure 7, we present intriguing findings regarding zero-shot generation outcomes.
Beyond modifying the camera’s extrinsic parameters, we also experimented with changing the intrinsic settings of the camera, configurations not encountered in the dataset. The successful generations demonstrate our model’s ability to generalize across different world volumes, indicating that it genuinely learns from world volume features instead of merely overfitting to the data.
We experimented with an unseen weather condition, "heavy snow," by altering the weather condition component of the scene guidance. This snow weather condition is entirely absent in the nuScenes dataset [2]. Nonetheless, our model incorporates the concept of snow from the Stable Diffusion pretrained model, integrating it into the scene based on its understanding of snow weather. This demonstrates the model’s ability to amalgamate concepts from the pretrained model with the fine-tuned street view images, showcasing its versatility and adaptability in generating content under novel conditions.
Interestingly, we altered the location word to the entirely different city of Shanghai, which presents significant differences from Singapore and Boston. This modification resulted in the generation of images that depict foggy-like weather and the architectural style characteristic of Shanghai’s highways. This ability to adapt the visual representation to a distinct geographical context further illustrates the model’s capacity to inherit and apply concepts from the Stable Diffusion pretraining. It showcases the model’s remarkable flexibility in synthesizing elements that reflect specific environmental and cultural attributes associated with various locations.
We also explored out-of-domain modifications in object guidance by substituting the word "car" with "purple car." Purple cars are quite rare in the dataset. Nonetheless, our model effectively integrated the concept of purple, successfully altering all cars in the scene to purple. This experiment underscores the model’s ability to comprehend and apply color attributes to specific objects within a scene, demonstrating its proficiency in adapting to and implementing novel characteristics not commonly found in the training data.
Figure 7: Examples of out-of-domain generation
0.A.3 Downstream tasks